OpenStudy (oswaldmurphy):

The distance traveled, in feet, of a ball dropped from a tall building is modeled by the equation d(t) = 16t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 2 to t = 5 represent?

OpenStudy (oswaldmurphy):

The ball travels an average distance of 112 feet from 2 seconds to 5 seconds. The ball falls down with an average speed of 48 feet per second from 2 seconds to 5 seconds. The ball falls down with an average speed of 112 feet per second from 2 seconds to 5 seconds. The ball travels an average distance of 48 feet from 2 seconds to 5 seconds. Please explain how to do this!

OpenStudy (welshfella):

average speed = distance interval / time interval

OpenStudy (welshfella):

t goes from 2 to 5 and you can work out the distances by plugging in t = 2 and t = 5 into d(t) = 16t^2

OpenStudy (welshfella):

in other words average speed is 16(5)^2 - 16(2)^2 --------------- 5 - 2

OpenStudy (oswaldmurphy):

Ok. Thank you!

OpenStudy (oswaldmurphy):

Is the answer A?

OpenStudy (welshfella):

400 - 64 / 3 = 336 / 3 = 112 feet per second

OpenStudy (welshfella):

112 is NOT a distance . 112 is a speed

OpenStudy (welshfella):

Not A.

OpenStudy (oswaldmurphy):

Oh I didnt notice that :P C.

OpenStudy (welshfella):

yes C is the correct option.